E. Mr. X is known to hit the target in 4 out of 5 shots, where as Mr. Y is known to hit the target in 3 out of 5 shots. Find the probability that both will hit the Target


1
Expert's answer
2022-01-14T06:49:16-0500

"p(X)={4\over5}" and "p(Y)={3\over5}"

Probability that both will hit the target is,

"p(hitting \space target)=p(X)\times p(Y)={4\over5}\times{3\over5}={12\over25}"

Therefore, the probability that both will hit the target is "{12\over25}".